OpenStudy (anonymous):

-2(a-8)+6a=36 what is the value of a?

OpenStudy (anonymous):

help! please

OpenStudy (midhun.madhu1987):

Applying Distributive Property, -2a + 8 + 6a = 36 -2a + 6a = 36 - 8 Find a.

OpenStudy (anonymous):

thank you!